### v1

# Regularized Projection Matrix Approximation with Applications to Community Detection

Abstract
This paper introduces a regularized projection matrix approximation framework aimed at recovering cluster information from the affinity matrix. The model is formulated as a projection approximation problem incorporating an entry-wise penalty function. We explore three distinct penalty functions addressing bounded, positive, and sparse scenarios, respectively, and derive the Alternating Direction Method of Multipliers (ADMM) algorithm to solve the problem. Then, we provide a theoretical analysis establishing the convergence properties of the proposed algorithm. Extensive numerical experiments on both synthetic and real-world datasets demonstrate that our regularized projection matrix approximation approach significantly outperforms state-of-the-art methods in terms of clustering performance.
Index Terms:
Projection Matrix Approximation, Regularisation, ADMM, Clustering
I Introduction
Community detection is a crucial problem in unsupervised learning that has garnered attention from researchers across various disciplines, including mathematics, statistics, physics, and social sciences. The objective is to partition $n$ data points into $K$ groups based on their pairwise similarities, represented by a similarity matrix $A\in{\mathbb{R}}^{n\texttimes n}$. A prevalent approach to tackle this problem involves first deriving a lower-dimensional representation of the data from $A$. Subsequently, a clustering algorithm such as $k$-means or the EM algorithm is applied to identify the clusters. The success of this method depends on the quality of the data representation and the accuracy of the computational methods used for $A$.
A popular approach for cluster identification is to utilize the top $K$ eigenvectors of matrix $A$, as employed in spectral clustering [1, 2]. Identifying these eigenvectors is equivalent, up to rotations, to determining the subspace spanned by them. This, in turn, is analogous to solving the following projection matrix approximation problem:
$\widehat{X}=\mathop{\mathrm{argmin}}_{X\in\mathcal{P}_{K}}\|A-X\|^{2}_{\rm F},$ (1)
where $\mathcal{P}_{K}\subseteq\mathbb{R}^{n\times n}$ is the set of rank-$K$ projection matrices. Thus, the effectiveness of spectral clustering is highly dependent on the quality of the projection matrix approximation.
Similar to the affinity matrix, the projection matrix also encapsulates class information. Consider an extreme scenario where clustering information is known beforehand, with $K$ classes, and each class consists of $\{n_{k},k=1,...,K\}$ samples. In this setup, the affinity matrix $A$ is defined such that $A_{i,j}=1$ if and only if samples $i$ and $j$ belong to the same group, and $A_{i,j}=0$ otherwise. Solving the corresponding eigenvalue problem provides the solution for (1) as follows:
$\widehat{X}_{i,j}=\left\{\begin{array}[]{cc}0,&A_{i,j}=0,\\ 1/n_{k},&A_{i,j}=1,\ c(i)=c(j)=k.\end{array}\right.$ (2)
where $c(\cdot)$ is the function which maps the sample index to its corresponding cluster number. It can be verified that $\widehat{X}$ is a low rank, non-negative, sparse projection matrix satisfying $\widehat{X}^{2}=\tilde{X}$. If we shuffle the indices and rearrange them according to the cluster indicator, we obtain a new projection matrix, which can be written as
$P\widehat{X}P^{T}=\left[\begin{array}[]{cccc}\frac{1}{n_{1}}{\bf 1}_{n_{1}\times n_{1}}&\bf 0&...&\bf 0\\ \bf 0&\frac{1}{n_{2}}{\bf 1}_{n_{2}\times n_{2}}&...&\bf 0\\ ...&...&...&...\\ \bf 0&...&...&\frac{\bf 1}{n_{K}}{\bf 1}_{n_{K}\times n_{K}}\end{array}\right],$
where ${\bf 1}_{n\times n}\in\mathbb{R}^{n\times n}$ is a squared matrix which all elements equal to $1$, and $P$ is a permutation matrix. We can observe that both $P\widehat{X}P^{T}$ and $\widehat{X}$ contain exactly of $\sum_{k}n_{k}^{2}$ nonzero elements.
In practical scenarios for community detection, the cluster information is typically not accessible beforehand. The affinity matrix $A$ is often computed using a kernel function or a cosine similarity function, which may deviate from the ideal assignment matrix. Persisting in applying the spectral projection approximation to derive the optimal rank-$K$ projection matrix approximation can result in a matrix $X$ with negative elements.
To address this issue, we propose leveraging additional information to regularize the projection approximation problem. This approach aims to obtain a higher-quality projection matrix $X$, thereby improving clustering outcomes. By incorporating regularization techniques, we can enforce non-negativity and sparsity, which align better with the true underlying cluster structure. This leads to more accurate and meaningful clustering results, even when the initial affinity matrix is imperfect.
Inspired by the optimal solution presented in equation (2), we propose that an effective projection matrix for clustering should exhibit certain distinctive properties such as non-negativity, boundedness, and a degree of sparsity. In this work, we introduce a uniform regularized projection problem approximation framework aimed at boosting clustering performance. Specifically, we study the entry-wise regularized projection approximation problem to restrict the entries of the projection matrix within a reasonable scope by considering:
$\min_{X\in\mathcal{P}_{K}\cap{\cal C}}\|A-X\|_{\rm F}^{2}.$ (3)
where $\mathcal{P}_{K}$ is the set consisting of all rank-$K$ projection matrices and $\cal C$ is some constraint set such as the pairwise bounded restriction $\{X|X_{i,j}\in[\alpha,\beta],\forall i,j\}$, the non-negative constraint $\{X|X_{i,j}\geq 0,\forall i,j\}$ or even the sparsity constraint such as $\{X|\|X\|_{0}\leq\eta\}$. Due to the simultaneously existence of the projection and extra constraint $\cal C$, it is challenging to directly solve (3). To address this difficulty, this paper proposes a regularized projection matrix approximation (RPMA) problem in (4) as an approximation to the above strict constrained problem, which is defined by
$\min_{X\in\mathcal{P}_{K}}\|A-X\|_{\rm F}^{2}+\lambda\sum_{ij}g(X_{ij}),\ \ g\in{\cal G}.$ (4)
where $\lambda$ serves as a non-negative tunable parameter, governing the trade-off between approximation error and entry-wise constraint. Additionally, $\cal G$ represents the class of positive, convex functions whose derivatives are continuous and satisfy the Lipschitz condition, i.e.,
${\cal G}:=\{g|g\in C^{(1)},$ $g\geq 0,g^{\prime\prime}(x)\geq 0,$
$|g^{\prime}(x_{1})-g^{\prime}(x_{2})|\leq\ell|x_{1}-x_{2}|,\forall x_{1},x_{2}\}.$
The construction of $\cal G$ is driven by two key considerations: Firstly, it must be expansive enough to adequately encapsulate our objectives as a penalty term. Secondly, it should adhere to specific properties that facilitate algorithm development and validate algorithm’s convergence behaviors.
Our contributions are threefold. First, we present a novel regularized projection matrix approximation framework designed to enhance the classical spectral clustering method. This framework could include three different types of penalty functions tailored for bounded, nonnegative, and sparse scenarios. Second, we develop an ADMM algorithm specifically for our model and rigorously analyze its convergence behavior. Third, we demonstrate the effectiveness of our approach through comprehensive validation using both synthetic and real-world datasets.
I-A Related works
Low-rank matrix optimization with additional structural constraints is a prevalent problem in machine learning and signal processing [3, 4]. The objective is to find the best low-rank matrix approximation that also satisfies certain structural constraints, such as non-negativity, symmetry, boundedness, simplex, and sparsity  [5, 6, 7]. There are two primary approaches to achieve this target. Firstly, these constraints can be enforced via matrix factorization with explicit constraints. Examples include non-negative matrix factorization [8], semi-nonnegative matrix factorization [9], bounded low-rank matrix approximation [10], and bounded projection matrix approximation [11]. Secondly, constraints can also be achieved via the soft-regularization term approach. This method includes techniques like simultaneously low-rank and sparse matrix approximation [12, 13, 14, 15]. However, these works primarily focus on seeking a low-rank matrix as a low-dimensional embedding of the input similarity matrix. In contrast, our work diverges from this goal. Instead of learning an embedding, we aim to recover a high-quality group connection matrix through regularized projection matrix approximation.
Motivated by the spectral clustering method and the unique attributes of the projection matrix $\widehat{X}$ in equation (2) (which represents the optimal approximation of the underlying true group connection $A$), we present a framework for regularized projection matrix approximation. This framework accommodates scenarios with bounded, non-negative, and sparse constraints. The varied regularization techniques aim to capture the structural intricacies of the true group connection from diverse perspectives.
II Constraints and Penalties
In this section, we study different forms of penalty function $g(\cdot)$ for achieving the various targeted constrains such as non-negativity, boundedness, and sparsity.
II-A Bounded Penalty
Based on (2), it’s evident that the low-rank approximation $X$ of an assignment matrix must be confined within $[0,1/\min_{k}n_{k}]$. Hence, it’s imperative for us to learn a bounded rank-$K$ projection matrix from $A$. In a general form, we address the problem as finding a projection matrix where each element falls within $[\alpha,\beta]$. One straightforward approach is to introduce the indicator function $I(x)$, defined such that $I(x)=0$ when $x$ lies in $[\alpha,\beta]$, and $I(x)=+\infty$ otherwise. However, this function’s lack of continuity poses difficulties when employing differentiable tools for analysis. To mitigate this, we introduce a smooth, convex function $g_{\alpha,\beta}(z)$ as a relaxation of the bounded indicator function, which is defined as
$g_{\alpha,\beta}(z)=(\min\{z-\alpha,0\})^{2}+(\min\{\beta-z,0\})^{2}.$ (5)
The construction of $g_{\alpha,\beta}(z)$ reveals that it solely penalizes $z$ when it lies outside the range $[\alpha,\beta]$. It can be observed that as $\lambda$ approaches +$\infty$+ ∞, $\lambda g_{\alpha,\beta}(z)$ exhibits behavior akin to that of the indicator function. Moreover, this function is convex, and its derivative is Lipschitz continuous, satisfying $|g^{\prime}_{\alpha,\beta}(x_{1})-g^{\prime}_{\alpha,\beta}(x_{2})|\leq 2|x_{1}-x_{2}|$.
II-B Non-negativity Penalty
The non-negativity constraint can be viewed as a one-side boundedness requirement by setting $\alpha=0,\beta=+\infty$. Thus, we can consider a non-negativity penalty function as a special case of the bounded penalty in (5). Therefore, we introduce this non-negative penalty function to regularize the projection matrix, which is defined as
$g(z):=(\min\{z,0\})^{2}.$ (6)
Similarly to the bounded penalty, the function $g(z)$ only exerts an effect when $z<0$. Furthermore, as $\lambda$ tends to infinity, $\lambda g(z)$ exhibits behavior akin to the indicator function $I(z)$, where $I(z)=0$ if and only if $z\geq 0$, and $I(z)=+\infty$ otherwise. Additionally, $g(z)$ is nonnegative and convex, and its derivative is Lipschitz continuous, satisfying $|g^{\prime}(x_{1})-g^{\prime}(x_{2})|=2|x_{1}-x_{2}|$.
II-C Sparsity Penalty
Figure 1: Illustration demonstrating why minimizing the $\ell_{1}$ norm leads to a sparse solution. The left diagram shows the curve $r(x_{1},x_{2})$ with $r=|x_{1}|+|x_{2}|$ and $x_{1}^{2}+x_{2}^{2}=1$, while the right diagram depicts the surface $r(x_{1},x_{2},x_{3})$ with $r=|x_{1}|+|x_{2}|+|x_{3}|$ and $x_{1}^{2}+x_{2}^{2}+x_{3}^{2}=1$.
The Lasso penalty [16] encourages sparsity by inducing certain coefficients or parameters in a model to become exactly zero. The rational for minimizing the $\ell_{1}$ norm yields a sparse solution can be illustrated by an example. In Figure 1, we provide a unit circle to represent the data $x\in\mathbb{R}^{2}$ (left diagram) and $x\in\mathbb{R}^{3}$ (right diagram) with fixed $\ell_{2}$ norm, and also plot the data by rescaling with the $\ell_{1}$ as $y=\|x\|_{1}x$. From Figure 1, we observe that $\min_{\|x\|_{2}=1}\|x\|_{1}$ achieves four optimal solutions: $(1,0),(-1,0),(0,1),(0,-1)$ when $x\in{\mathbb{R}}^{2}$ in the left diagram, and six optimal solutions: $(1,0,0),(-1,0,0),(0,1,0),(0,-1,0),(0,0,1)$ and $(0,0,-1)$ when $x\in{\mathbb{R}}^{3}$ in the right diagram. Since any rank-$K$ projection matrix has a fixed Frobenius norm equal to $\sqrt{K}$, we speculate that $\min_{X\in{\cal P}_{K}}\|X\|_{1}$ will also yield a sparse solution, similar to the examples in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$.
Motivated by the example in Figure 1, it is natural to set $g(z)$ as the absolute value function $|z|$ to regularize the projection approximation problem. Consequently, the summation of $|X_{i,j}|$ leads to the Lasso penalty function $\|X\|_{1}$ such that $\sum_{i,j}g(X_{i,j})=\|X\|_{1}$. However, the derivative of the absolute value function is non-differentiable at 00 and does not satisfy the Lipschitz continuous property. This can be verified as follows: for any $x_{1}\rightarrow 0^{-},x_{2}\rightarrow 0^{+}$, we have $|g^{\prime}(x_{1})-g^{\prime}(x_{2})|=2$, while $|x_{1}-x_{2}|$ can be arbitrarily small. Therefore, the absolute function does not belong to $\cal G$.
To address this, we can approximate of the absolute function by introducing the Huber loss function [17, 18], which smooths the absolute value function with a quadratic function in the neighborhood of 00. The Huber loss function is defined by
$g_{\delta}(x)=\left\{\begin{array}[]{ll}\frac{x^{2}}{2\delta},&|x|\leq\delta,\\ |x|-\frac{\delta}{2},&{\rm otherwise},\end{array}\right.$ (7)
where $\delta$ is a positive threshold parameter. It can be verified that Huber function retains the advantages of the absolute value function while ensuring both differentiability and Lipschitz continuity at $x=0$.
It is worth mentioning that as $\delta\rightarrow 0^{+}$, $g_{\delta}(x)$ uniformly converges to $|x|$. Thus, the Huber loss function can induce sparse solutions similarly to the Lasso penalty when $\delta$ is small enough. We propose using the summation of the Huber loss function $\sum_{i,j}g_{\delta}(X_{i,j})$ as a substitute for the Lasso penalty $\|X\|_{1}$ to learn a sparse projection approximation by adopting a sufficient small $\delta$. Compared to the absolute value function, the Huber loss function possesses specific advantages, such as smoothness and Lipschitz continuity, demonstrated by $|g^{\prime}_{\delta}(x_{1})-g^{\prime}_{\delta}(x_{2})|\leq\frac{1}{\delta}|x_{1}-x_{2}|.$
III Solution to Univariate Problems
The key strategy for tackling (4) revolves around segregating the constraint $X\in{\cal P}_{K}$ from the objective function by introducing a new variable $Y$. This transformation is executed albeit at the expense of appending two constraints: $X=Y$ and $Y\in{\cal P}_{K}$. Noticing that the squared Frobenius norm is a separable function, we first drop the constraint $X\in{\cal P}_{K}$ and rewrite (4) as a problem
$\min_{\{X_{i,j}\}}\sum_{i,j}\{(A_{i,j}-X_{i,j})^{2}+\lambda g(X_{i,j})\}.$ (8)
The optimization problem with respect to variables $\{X_{i,j}\}$ are separable in (8), allowing the optimal minimization to be decomposed into smaller, more manageable subproblems. Consequently, it is essential to investigate the unconstrained univariate problem, formulated as follows:
$f(z)=(s-z)^{2}+\tau g(z).$
where $\tau>0$, and $g(z)$ represents a convex, differentiable function, which can take various forms such as those depicted in (5), (6), and (7). The convexity of $g(z)$ ensures that $\min_{z}f(z)$ possesses a unique optimal solution. Unraveling this solution not only yields the optimal outcome for $\min_{z}f(z)$ but also offers valuable insights into our regularized projection approximation problem.
III-A Bounded Penalty
For the bounded loss scenario, we set $g(z)$ as $g(z):=g_{\alpha,\beta}(z)$, where $\alpha<\beta$, as specified in (5). Subsequently, we proceed to derive the closed-form solution for $\min_{z\in\mathbb{R}}(s-z)^{2}+\tau g_{\alpha,\beta}(z)$. We rewrite this problem as
$\min_{z\in\mathbb{R}}\big{\{}(s-z)^{2}+\tau\{(\min\{z-\alpha,0\})^{2}+(\min\{\beta-z,0\})^{2}\}\big{\}}.$ (9)
Depending on the magnitudes of $z$, $\alpha$, and $\beta$, the objective function takes on three distinct forms across three specific intervals. We will analyze and derive the optimal solution for each interval individually to gain a comprehensive understanding of the overall minimization problem.
Firstly, let us examine the problem (9) restricted to the interval $z\leq\alpha<\beta$. In this case, (9) simplifies to:
$\widehat{z}_{1}=\arg\min_{z\leq\alpha}(z-s)^{2}+\tau(z-\alpha)^{2}.$
Therefore, for this sub-problem, the optimal solution is attained by selecting the smaller value between $\alpha$ and $\frac{s+\tau\alpha}{\tau+1}$, resulting in: $\widehat{z}_{1}=\min\{\alpha,\frac{s+\tau\alpha}{\tau+1}\}=\frac{\tau\alpha+\min\{\alpha,s\}}{\tau+1}$.
Secondly, let us examine the sub-problem by constraining $z\geq\beta$. In this scenario, the problem (9) over this interval simplifies to:
$\widehat{z}_{2}=\arg\min_{z\geq\beta}(z-s)^{2}+\tau(\beta-z)^{2}.$
Similarly, the optimal solution for the problem constrained by $z\geq\beta$ is achieved by selecting the larger value between $\beta$ and $\frac{s+\tau\beta}{\tau+1}$, i.e., $\widehat{z}_{2}=\max\{\beta,\frac{s+\tau\beta}{\tau+1}\}=\frac{\tau\beta+\max\{\beta,s\}}{\tau+1}$.
Thirdly, let us consider the case where $\alpha\leq z\leq\beta$. In this scenario, the problem (9) simplifies to:
$\widehat{z}_{3}=\arg\min_{\alpha\leq z\leq\beta}(z-s)^{2}.$
Similarly to the previous cases, the optimal solution for the problem constrained by $\alpha\leq z\leq\beta$ is given by: $\widehat{z}_{3}=\max\{\alpha,\min\{\beta,s\}\}$.
TABLE I: The values of $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ along with their respective function values $f(\widehat{z}_{1}),f(\widehat{z}_{2}),f(\widehat{z}_{3})$ are provided for three distinct magnitude scenarios of $\alpha,\beta$, and $s$.
$\widehat{z}_{1}$ $\widehat{z}_{2}$ $\widehat{z}_{3}$ $f(\widehat{z}_{1})$ $f(\widehat{z}_{2})$ $f(\widehat{z}_{3})$ $\arg\min_{k}f(\widehat{z}_{k})$
$s\leq\alpha$ $\frac{\tau\alpha+s}{\tau+1}$ $\beta$ $\alpha$ $\frac{\tau(\alpha-s)^{2}}{\tau+1}$ $(\beta-s)^{2}$ $(\alpha-s)^{2}$ $\frac{\tau\alpha+s}{\tau+1}$
$\alpha\leq s\leq\beta$ $\alpha$ $\beta$ $s$ $(\alpha-s)^{2}$ $(\beta-s)^{2}$ 00 $s$
$\beta\leq s$ $\alpha$ $\frac{\tau\beta+s}{\tau+1}$ $\beta$ $(\alpha-s)^{2}$ $\frac{\tau(\beta-s)^{2}}{\tau+1}$ $(\beta-s)^{2}$ $\frac{\tau\beta+s}{\tau+1}$
Finally, the global optimal solutions can be determined by comparing the optimal function values across each interval. Given that $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ are the optimal solutions for (9) on three different intervals, it is necessary to evaluate the function values at $\widehat{z}_{1},\widehat{z}_{2}$, and $\widehat{z}_{3}$. Table I presents the specific values of $\widehat{z}_{k},f(\widehat{z}_{k})$ for $k=1,2,3$, covering three distinct magnitude scenarios of $s$, $\alpha$, and $\beta$. By analyzing the magnitudes of $f(\widehat{z}_{1}),f(\widehat{z}_{2}),f(\widehat{z}_{3})$, we can determine that the global optimal solution is:
$\arg\min_{z}f(z)=\left\{\begin{array}[]{cc}\frac{\tau\alpha+s}{\tau+1},&s\leq\alpha,\\ s,&\alpha\leq s\leq\beta,\\ \frac{\tau\beta+s}{\tau+1},&\beta\leq s.\end{array}\right.$ (10)
In a compact form, the solution $\widehat{z}$ can be written as $\widehat{z}=\frac{1}{1+\tau}(s+\tau\mathcal{P}_{\alpha,\beta}(s)),$ where $\mathcal{P}_{\alpha,\beta}(\cdot)$ is an projection operator given by $\mathcal{P}_{\alpha,\beta}(s)=\min\{\max\{s,\alpha\},\beta\}$. Referring to equation (10), it’s evident that for any value of $s$ outside the range $[\alpha,\beta]$, the optimal solution can be expressed as a convex combination $ws+(1-w)\mu$, where $w=\frac{1}{\tau+1}$ and $\mu=\arg\min_{t\in[\alpha,\beta]}|t-s|$. As $\tau$ tends to infinity, we obtain $\widehat{z}\rightarrow\mathcal{P}_{\alpha,\beta}(s)$.
III-B Non-negativity Penalty
Similarly to section III-A, we set the penalty term for the non-negative scenario as $g(z):=(\min\{z,0\})^{2}$. This penalty term penalizes negative values of $z$, encouraging non-negativity. Therefore, we need to derive the closed-form solution to the following optimization problem:
$\min_{z}\{(s-z)^{2}+\tau(\min\{z,0\})^{2}\}.$
To solve this, we consider two cases based on the sign of $z$:
Case 1: For $z\leq 0$, we have $\min\{z,0\}=z$, the problem simplifies to:
$\widehat{z}_{1}=\arg\min_{z\leq 0}(1+\tau)z^{2}-2sz+s^{2}=\min\{0,s/(1+\tau)\}.$
Case 2: For $z\geq 0$, we have $\min\{z,0\}=0$, the problem simplifies to:
$\widehat{z}_{2}=\arg\min_{z\geq 0}(z-s)^{2}=\max\{0,s\}.$
Finally, we merge the results of two cases and locate the global optimal solution. This can be achieved by comparing the function value $f(\cdot)$ at $\widehat{z}_{1}$ and $\widehat{z}_{2}$. Since $\widehat{z}_{1}$ and $\widehat{z}_{2}$ depends on $s$, to compare the magnitude of $f(\widehat{z}_{1})$ and $f(\widehat{z}_{2})$, we need to separate them into two cases depending on the sign of $s$:
Case 1: For $s\leq 0$, we have $\widehat{z}_{1}=s/(1+\tau)$ and $\widehat{z}_{2}=0$. Thus, $f(\widehat{z}_{1})=\frac{s^{2}\tau}{1+\tau}<s^{2}=f(\hat{z}_{2})$. Since $\tau>0$, the optimal solution is $\widehat{z}=\frac{1}{1+\tau}$.
Case 2: For $s>0$, we have $\widehat{z}_{1}=0$ and $\widehat{z}_{2}=s$. Here, $f(0)=s^{2}>f(s)=0$. Therefore, the optimal solution is $\widehat{z}=s$.
To summarize, the optimal solution is given by:
$\arg\min_{z}f(z)=\left\{\begin{array}[]{lc}\frac{s}{1+\tau},&s\leq 0,\\ s,&s>0.\end{array}\right.$ (11)
It can be observed from (11) that the solution reduces the scale for non-negative $s$ by a factor of $\frac{1}{1+\tau}$, while leaving positive $s$ unchanged.
III-C Sparsity Penalty
For the Huber loss penalty, we set $f_{\delta}(z):=(z-s)^{2}+\tau g_{\delta}(z)$, where $\delta$ is the hyper-parameter determines the transition point between the quadratic function and the absolute value components of the Huber loss. In what follows, we investigate how the Huber penalty $g_{\delta}(z)$ influences the solution of $f_{\delta}(z)$ by discussing three cases depending on the magnitude of $z$ relative to $\delta$.
Case 1: For $|z|\leq\delta$, the optimization problem $\min_{|z|\leq\delta}f_{\delta}(z)$ simplifies into:
$\widehat{z}_{1}$ $=\arg\min_{|z|\leq\delta}\{(1+\frac{\tau}{2\delta})z^{2}-2zs+s^{2}\}$
$=\arg\min_{|z|\leq\delta}(1+\frac{\tau}{2\delta})(z-\frac{2\delta s}{2\delta+\tau})^{2},$
The optimal solution can be expressed as $\widehat{z}_{1}=\max\{-\delta,\min\{\frac{2\delta s}{2\delta+\tau},\delta\}\}$, which can be viewed as a projection of $\frac{2\delta s}{2\delta+\tau}$ onto the interval $[-\delta,\delta]$.
Case 2: For $z\geq\delta$, the optimization problem $\min_{z\geq\delta}f_{\delta}(z)$ simplifies into
$\widehat{z}_{2}=\arg\min_{z\geq\delta}(z-s)^{2}+\tau z=\arg\min_{z\geq\delta}(z-(s-\frac{\tau}{2}))^{2}+s^{2}.$
Similarly, the optimal solution yields the closed-form as $\widehat{z}_{2}:=\max\{\delta,s-\frac{\tau}{2}\}$, which is the projection of $s-\frac{\tau}{2}$ onto $[\delta,+\infty)$
Case 3: For $z\leq-\delta$, the objective function simplifies into:
$\widehat{z}_{3}=\arg\min_{z\leq-\delta}(z-s)^{2}-\tau z=\arg\min_{z\leq-\delta}(z-(s+\frac{\tau}{2}))^{2}+s^{2}.$
Similarly to the above two cases, this problem yields the closed-form solution as $\widehat{z}_{3}:=\min\{-\delta,s+\frac{\tau}{2}\}$ and it can be viewed as the projection of $s+\frac{\tau}{2}$ onto $(-\infty,-\delta]$.
The optimal solution of $f_{\delta}(z)$ corresponds to the optimal solutions at $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ over the intervals $|z|\leq\delta$, $z\geq\delta$, and $z\leq-\delta$, respectively. Due to the intricate relationships among $f(\widehat{z}_{1}),f(\widehat{z}_{2}),f(\widehat{z}_{3})$, explicitly determining the smallest value becomes challenging. Nonetheless, in practical scenarios, we can still numerically identify the global solution of $f_{\delta}(z)$ by evaluating the function $f_{\delta}(\cdot)$ at $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ and selecting the one with the smallest function value as the optimal solution to $\min_{z}f_{\delta}(z)$, i.e.,
$\widehat{z}=\arg\min_{k=1,2,3}f_{\delta}(\widehat{z}_{k}).$ (12)
Table II illustrates the representation of $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$, along with the corresponding function values $f_{\delta}(\widehat{z}_{1}),f_{\delta}(\widehat{z}_{2}),f_{\delta}(\widehat{z}_{3})$, contingent upon the magnitude relationship among $s,\tau$ and $\delta$. Therefore, for any given $s,\tau$ and $\delta$, we can get the optimal value of $f_{\delta}(\cdot)$ by comparing $f_{\delta}(\widehat{z}_{1}),f_{\delta}(\widehat{z}_{2})$, $f_{\delta}(\widehat{z}_{3})$. We then set the optimal solution to $f_{\delta}(z)$ as the value from $\{\widehat{z}_{k},k=1,2,3\}$ that achieves the smallest value of $f_{\delta}(\widehat{z}_{k}),k=1,2,3$. The optimal value of $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ and their respective function values are listed in Table II.
We have now concluded the discussion on the univariate problem. Next, we will explore the solution of the regularized projection matrix approximation in (4) using the alternating direction method of multipliers (ADMM) [19, 20].
TABLE II: The values of $\widehat{z}_{1},\widehat{z}_{2},\widehat{z}_{3}$ along with their respective function values $f(\widehat{z}_{1}),f(\widehat{z}_{2}),f(\widehat{z}_{3})$ are provided for three distinct magnitude scenarios of $\alpha,\beta$, and $s$.
$s+\frac{\tau}{2}\leq-\delta$ $s-\frac{\tau}{2}\leq-\delta\leq s+\frac{\tau}{2}$ $-\delta\leq s-\frac{\tau}{2}\leq s+\frac{\tau}{2}\leq\delta$ $s-\frac{\tau}{2}\leq\delta\leq s+\frac{\tau}{2}$ $\delta\leq s-\frac{\tau}{2}$
$\widehat{z}_{1}$ $-\delta$ $\frac{2\delta s}{2\delta+\tau}$ $\frac{2\delta s}{2\delta+\tau}$ $\frac{2\delta s}{2\delta+\tau}$ $\delta$
$\widehat{z}_{2}$ $\delta$ $\delta$ $\delta$ $\delta$ $s-\frac{\tau}{2}$
$\widehat{z}_{3}$ $s+\frac{\tau}{2}$ $-\delta$ $-\delta$ $-\delta$ $-\delta$
$f(\widehat{z}_{1})$ $(\delta+s)^{2}+\frac{\tau\delta}{2}$ $\frac{\tau s^{2}}{2\delta+\tau}$ $\frac{\tau s^{2}}{2\delta+\tau}$ $\frac{\tau s^{2}}{2\delta+\tau}$ $(\delta-s)^{2}+\frac{\tau\delta}{2}$
$f(\widehat{z}_{2})$ $(\delta-s)^{2}+\tau\delta$ $(\delta-s)^{2}+\tau\delta$ $(\delta-s)^{2}+\tau\delta$ $(\delta-s)^{2}+\tau\delta$ $s^{2}$
$f(\widehat{z}_{3})$ $s^{2}$ $(\delta+s)^{2}+\tau\delta$ $(\delta+s)^{2}+\tau\delta$ $(\delta+s)^{2}+\tau\delta$ $(\delta+s)^{2}+\tau\delta$
IV Algorithm
In this section, we develop an ADMM algorithm [19, 20] to solve the RPMA framework. First, we define the augmented Lagrangian ${\cal L}_{\rho}(X,Y,\Lambda)$ as
${\cal L}_{\rho}(X,Y,\Lambda)=\|A-X\|_{\rm F}^{2}+$ $\lambda\sum_{ij}g(Y_{ij})$
$+\frac{\rho}{2}\|X-Y\|_{\rm F}^{2}$ $+\langle\Lambda,X-Y\rangle,$
where $g(\cdot)\in\cal G$ can be chosen to suit various regularization purposes, such as imposing bounded, non-negative, or sparse penalties, as discussed in this work.
Maximizing ${\cal L}_{\rho}(X,Y,\Lambda)$ with respect to the dual variable $\Lambda$ necessitates $X=Y$, otherwise it diverges to infinity. This implicit constraint $X=Y$ renders our original problem (4) equivalent to the following optimization problem:
$\min_{Y,X\in{\cal P}_{K}}\max_{\Lambda}{\cal L}_{\rho}(X,Y,\Lambda)$ (13)
The alternating direction method of multipliers solves (13) by constructing an iterative sequence as follows: Starting from initialization points $\{X^{0},Y^{0},\Lambda^{0}\}$, our algorithm updates $\{X^{k},Y^{k},\Lambda^{k}\}$ alternatively as:
$X^{k+1}$ $=\mathop{\mathrm{argmin}}_{X\in{\cal P}_{K}}{\cal L}_{\rho}(X,Y^{k},\Lambda^{k}),$ (14)
$Y^{k+1}$ $=\mathop{\mathrm{argmin}}_{Y}{\cal L}_{\rho}(X^{k+1},Y,\Lambda^{k}),$ (15)
$\Lambda^{k+1}$ $=\Lambda^{k}+\rho(X^{k+1}-Y^{k+1}).$ (16)
These updates in (14) and (15) have closed-form solutions and thus can be implemented efficiently. Specifically, problem (14) is equivalent to the following problem:
$X^{k+1}=\mathop{\mathrm{argmax}}_{X\in{\cal P}_{K}}\langle X,W^{k}\rangle,\quad W^{k}=2A+\rho Y^{k}-\Lambda^{k}.$
This problem of minimizing the inner product of the projection matrix can be effectively addressed by utilizing the eigenvalue decomposition of $W^{k}$ as follows:
$\widehat{U}=\max_{U^{T}U=I}{\rm trace}(U^{T}W^{k}U).$
Therefore, $X^{k+1}$ is given by the projection matrix associated with the leading $K$ eigenvectors of $W^{k}$, i.e., $X^{k+1}=\widehat{U}\widehat{U}^{T}$.
The update with respect to $Y$ can be reformulated as an equivalent problem (15) as:
$Y^{k+1}=\mathop{\mathrm{argmin}}_{Y}\|Y-V^{k+1}\|_{\rm F}^{2}+\tau\sum_{ij}g(Y_{ij}),$
where $V^{k+1}=X^{k+1}+\Lambda^{k}/\rho$ and $\tau={2\lambda}/{\rho}$. This is a separable problem, and each entry $Y^{k+1}_{ij}$ can be solved via
$Y^{k+1}_{ij}=\mathop{\mathrm{argmin}}_{Y_{ij}}\left(Y_{ij}-V^{k+1}_{ij}\right)^{2}+\tau g(Y_{ij}).$ (17)
The closed-form solution of (17) varies depending on different motivations and form of $g(\cdot)$, such as boundedness, non-negativity, or sparsity regularization terms. These different penalty functions lead to different form of solution to (17), as seen in (10), (11), and (12).
IV-A Parameter Setting
Figure 2: The solutions of the sparse regularization model vary with different selections of $\lambda$ for a simulated dataset where $P(A_{i,j}=1)=0.65$ when $c(i)=c(j)$ and $P(A_{i,j}=1)=0.40$ when $c(i)\neq c(j)$.
In this section, we demonstrate how to set the parameter $\lambda$ for a specific regularization term. The appropriate choice of the trade-off parameter $\lambda$ depends on the formulation of the regularization terms. For example, when dealing with bounded and nonnegative penalties, as described in equations (5) and (6), where entries $X_{i,j}$ fall within $[\alpha,\beta]$ or $[0,+\infty)$ respectively, $\lambda$ can be chosen to be as large as possible to approximate the indicator function effectively.
Conversely, when utilizing the Huber loss penalty to promote sparsity within the projection matrix, the penalty loss function does not have the property of refraining from penalizing entries within a specified range, unlike the other two penalties. Therefore, we advocate for a more cautious selection of $\lambda$ to avoid generating excessively sparse solutions. In Figure 2, we provide a demonstration illustrating that a large $\lambda$ could result in an overly sparse solution.
V Convergence Theory
This section provides convergence properties of the proposed ADMM algorithm. We show that any limiting point of the solution sequence is a stationary point of problem (4) as long as $g(\cdot)\in\cal G$. Our proof consists of three components.
First, we establish Lemma 1, that the successive variation of the dual variable $\Lambda$ is determined by the corresponding change in $Y$, proportionally scaled by a constant factor of $2\lambda\ell$. Here, $\lambda$ and $\ell$ denote the trade-off parameter within the objective function and the Lipschitz parameter of $g(\cdot)$, respectively.
Lemma 1.
$\|\Lambda^{k+1}-\Lambda^{k}\|_{\rm F}\leq 2\lambda\ell\|Y^{k+1}-Y^{k}\|_{\rm F}$.
We postpone the proof of this lemma into the appendix. The validity of this inequality becomes apparent upon observing that $\nabla_{Y}{\cal L}(X^{k+1},Y,\Lambda^{k})|_{Y=Y^{k+1}}=0$ and considering the update of the dual variable in (16). Next, we show that ${\cal L}_{\rho}(X^{k},Y^{k},\Lambda^{k})$ is decreasing in $k$ and the difference is lower bounded by $\frac{\sqrt{3}}{6}\ell\lambda\|Y^{k+1}-Y^{k}\|_{\rm F}^{2}$ when we set $\rho=\sqrt{3}\lambda\ell$.
Lemma 2.
Let $\rho=\sqrt{3}\lambda\ell$. The following inequality holds:
$D\coloneqq{\cal L}_{\rho}(X^{k},Y^{k},\Lambda^{k})-{\cal L}_{\rho}(X^{k+1},Y^{k+1},\Lambda^{k+1})$
$\geq\frac{\sqrt{3}}{6}\ell\lambda\|Y^{k+1}-Y^{k}\|_{\rm F}^{2}.$ (18)
The proof of Lemma 2 is left in the appendix. The inequality in (18) can be verified by applying the convexity of $g(\cdot)$ and Lemma 1. Finally, we demonstrate that any limiting point of $\{X^{k},Y^{k},\Lambda^{k}\}$ is a stationary point of problem (4).
Theorem 1.
Let $\rho=\sqrt{3}\lambda\ell$ and $\{X^{*},Y^{*},\Lambda^{*}\}$ be a limiting point of $\{X^{k},Y^{k},\Lambda^{k}\}$. Suppose $W^{*}=2A+\rho Y^{*}-\Lambda^{*}$ has distinct $K$-th and $(K+1)$-th eigenvalues. Then, the Karush–Kuhn–Tucker (KKT) condition of problem (4) holds for the limiting point, i.e.,
{$&X^{*}=Y^{*},\ X^{*}\textnormal{ is the projection matrix associated}\\ &\quad\textnormal{ with the top }K\textnormal{ eigenvectors of }W^{*},\\ &\Lambda_{ij}^{*}+\rho(Y^{*}_{ij}-X_{ij}^{*})=\lambda g_{\alpha,\beta}^{\prime}(Y^{*}_{ij}),\quad\forall i,j,\end{aligned}\right.$ (19)
The proof of this theorem is postponed to the appendix. The theorem can be proved by applying the KKT condition [21] in the updating with respect to $X$ and $Y$ in (14) and (15), the non-negativity $g(z)$ together with the conclusion in Lemma 2.
VI Numerical Experiment
In this section, we begin with a synthetic example to illustrate the advantages of our regularized projection approximation problem. Following this, we compare the proposed RPMA with other community detection algorithms, including SDP-1 [22], SDP-2 [23], SLSA [12], and spectral clustering (SP)[2], using real-world datasets. The objective functions for the compared methods are summarized in TableIV. The real-world datasets used in this work include the Handwritten Digit dataset [24] , the Columbia University Image Library dataset [25]and the Human Activity Recognition dataset [26].
We give an introduction to the datasets used in our work. Firstly, the Handwritten Digit dataset is sourced from the UCI Machine Learning Repository [^1] https://archive.ics.uci.edu/ml/datasets/Multiple+Features, encompassing 10 classes ranging from 0 to 9, with each class containing 200 images. Each image consists of 240 pixels ($16\times 15$). We utilize the first 1000 samples from classes 0 to 4, naming this subset DIGIT5. Additionally, we employ all 2000 samples, referring to this comprehensive set as DIGIT10. Secondly, the Columbia University Image Library (COIL) is a dataset used in computer vision and image processing research. Created by the Center for Research in Intelligent Systems at Columbia University [^2] https://www.kaggle.com/datasets/jessicali9530/coil100, the COIL dataset contains gray scale images of 20 different objects. Each object was placed on a motorized turntable and imaged at pose intervals of 5 degrees, resulting in 72 images per object, making a total of 1,440 images in the dataset. The objects include a variety of common items such as toys, tools, and household objects. We name the images from the first 10 classes as COIL10 and all the images as COIL20. Finally, the Human Activity Recognition (HAR) dataset is also sourced from the UCI Machine Learning [^3] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones. It is a comprehensive dataset collected from 30 volunteers wearing smartphones with inertial sensors during various daily activities. The data has been pre-processed to reduce noise and extract features, which enables the classification of six distinct activities using acceleration and angular velocity measurements. We assign six labels (0 to 5) to the activity categories, draw 500 samples ranging from 0 to 2 and designate them as HAR3, and draw 1000 samples from the entire dataset, called HAR6.
VI-A Experiment with Synthetic Data
Figure 3: The illustration depicts the solutions corresponding to RPMA-Sparse, RPMA-Bound, and RPMA-Positive with different settings of $\lambda$.
TABLE III: Comparison of the performance of bounded, positive and
sparse regularized projection matrix approximation under different setting of parameters.
Frobenius norm: $\|P_{A}-P^{*}\|_{\rm F}$ ACC NMI
$\delta$ $\lambda=10$ $\lambda=10^{2}$ $\lambda=10^{3}$ $\lambda=10^{4}$ $\lambda=10^{5}$ $\lambda=10$ $\lambda=10^{2}$ $\lambda=10^{3}$ $\lambda=10^{4}$ $\lambda=10^{5}$ $\lambda=10$ $\lambda=10^{2}$ $\lambda=10^{3}$ $\lambda=10^{4}$ $\lambda=10^{5}$
Bounded 0.6938 0.8340 0.8531 0.8531 0.8523 0.9100 0.8750 0.9000 0.9000 0.9000 0.5670 0.4592 0.5310 0.5310 0.5310
Positive 0.7852 0.7621 1.0080 1.0080 1.0092 0.8750 0.9000 0.8500 0.8500 0.8250 0.4592 0.5310 0.3988 0.3988 0.3464
Sparse $\delta$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=0.3$ $\lambda=0.4$ $\lambda=0.5$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=0.3$ $\lambda=0.4$ $\lambda=0.5$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=0.3$ $\lambda=0.4$ $\lambda=0.5$
$10^{-3}$ 0.8713 0.9156 0.9084 0.9458 0.7307 0.8250 0.8313 0.8750 0.8750 0.9500 0.3464 0.3470 0.4592 0.4592 0.7136
$10^{-4}$ 0.8791 0.9386 0.9026 0.9421 0.7278 0.8250 0.8000 0.8750 0.8750 0.9500 0.3464 0.2832 0.4592 0.4592 0.7136
$10^{-5}$ 0.8515 0.8965 0.9025 0.9419 0.7276 0.8638 0.8500 0.8750 0.8750 0.9500 0.4320 0.3902 0.4592 0.4592 0.7136
$10^{-6}$ 0.8500 0.8934 0.9025 0.9419 0.7275 0.8650 0.8500 0.8750 0.8750 0.9500 0.4350 0.3902 0.4592 0.4592 0.7136
$10^{-7}$ 0.8498 0.8931 0.9025 0.9419 0.7275 0.8600 0.8500 0.8750 0.8750 0.9500 0.4229 0.3902 0.4592 0.4592 0.7136
TABLE IV: The objectives and constraints for related algorithms
SDP-1 SDP-2 Spectral SLSA
[…]
TABLE V: Performance Comparison of the Related Community Detection Algorithms
Criterion ACC NMI
Methods SDP-1 SDP-2 Spectral SLSA RPMA SDP-1 SDP-2 Spectral SLSA RPMA
Sparse Bound Positive Sparse Bound Positive
DIGIT5 0.922 0.919 0.929 0.928 0.952 0.938 0.924 0.818 0.816 0.819 0.818 0.877 0.832 0.813
DIGIT10 0.677 0.687 0.684 0.680 0.817 0.634 0.646 0.635 0.644 0.643 0.620 0.751 0.556 0.588
COIL10 0.541 0.550 0.550 0.546 0.602 0.560 0.556 0.596 0.605 0.605 0.570 0.654 0.576 0.604
COIL20 0.644 0.625 0.635 0.569 0.709 0.565 0.615 0.768 0.762 0.764 0.697 0.806 0.677 0.738
HAR3 0.640 0.638 0.643 0.642 0.701 0.664 0.651 0.606 0.605 0.612 0.554 0.640 0.591 0.586
HAR6 0.617 0.622 0.636 0.593 0.685 0.621 0.633 0.609 0.611 0.614 0.515 0.631 0.572 0.603
We generate a random symmetric community connection matrix $A$ such that $P(A_{i,j}=1)=p_{1},c(i)=c(j)$ and $P(A_{i,j}=0)=p_{2},c(i)\neq c(j)$. Then, we apply the projection matrix approach with the bounded, positive and sparse regularizations to denoise $A$ and obtained a regularized projection matrix $P_{A}$. By rearranging the elements using the clustering information, we know $\mathbb{E}(A)$ is a diagonal block matrix up to a permutation matrix $Q$. We derive the ideal rank-$K$ projection matrix $P_{*}$ via solving $\min_{X\in\cal P_{K}}\|{\mathbb{E}}(A)-X\|_{\rm F}^{2}$ and measure distance between the regularized solution $P_{A}$ and $P_{*}$. We apply $k$-means clustering on the rows the orthonormal matrix $U\in{\mathbb{R}}^{n\times K}$, where $U$ is the factorization of $P_{A}$ such that $P_{A}=UU^{T},U^{T}U=I_{K}$. The clustering performance is evaluated using two standard criteria: accuracy (ACC) and normalized mutual information (NMI) [27].
We set $p_{1}=0.65,p_{2}=0.40$ and the number of samples $n=40$ and number of community $K=2$. We set the parameters $\alpha=0,\beta=1/20$ for the bounded regularization approach. We list the result measured by Frobenius norm, ACC, and NMI in the Table III. The spectral clustering method achieves result as ACC=$0.815$, NMI = $0.328$, $\|P_{A}-P_{*}\|_{\rm F}=0.898$. In Figure 3, we plot the solutions $P_{A}$ corresponding to the bounded, positive and sparse regularization, and illustrate the rational behind in regularized projection approximation model.
We can conclude from Table III that: Firstly, utilizing the sparse regularization method enhances clustering performance significantly, elevating ACC from 0.815 to 0.950 and NMI from 0.328 to 0.714. Secondly, the bounded regularization approach achieves the closest approximation to the ideal projection matrix according to the Frobenius norm criterion, with $\|P_{A}-P_{*}\|_{\rm F}=0.6938$ when $\lambda=10$. Thirdly, each of the three regularization approaches proposed in this study contributes to enhancing clustering performance by leveraging additional structural information encoded in the projection matrix corresponding to the assignment matrix.
VI-B Experiment on Real-world Dataset
In this section, we demonstrate how RPMA can be used to enhance spectral clustering. We compare the results of RPMA with those of the semidefinite positive approximation and SLSA on the Coil, handwritten digit, and Human Activity Recognition datasets. The similarity matrix $A$ is constructed using the Gaussian kernel such that $A_{i,j}=\exp(-\|x_{i}-x_{j}\|_{2}^{2}/\sigma^{2})$ where the bandwidth parameter is set as $\sigma^{2}=\frac{2}{n(n-1)}\sum_{i<j}\|x_{i}-x_{j}\|_{2}^{2}$. We set the parameters $\delta\in\{10^{-3},10^{-4},10^{-5},10^{-6}\}$ and $\lambda=0.1\times k,k=1,2...,8$ in the sparse regularized projection approximation model and $\alpha=0,\beta=1/\min_{k}n_{k}$ for the bounded regularized projection approximation model. We use the result from spectral clustering as the initialization for all three approaches (sparse, bounded, positive) regularized projection matrix approximation problems.
Figure 4: Illustration of solutions corresponding to SDP-1, SDP-2, Spectral Clustering, SLSA, and RPMA Sparse for the COIL10 (top row) and DIGIT10 (bottom row) datasets, respectively.
We present the optimal performance in Table V and visualize the solutions of each algorithm in Figure 4. Upon comparing the solutions of SLSA and PRMA-Sparse, we observe that PRMA-Sparse effectively recovers the signal in each diagonal block, while SLSA exhibits significant signal loss, particularly in the 9th diagonal block. Furthermore, we deduce that RPMA-Sparse surpasses the other algorithms considered, as evidenced by achieving the highest ACC and NMI scores in Table V.
VII Conclusion
Spectral clustering is a widely used approach for community detection from a connection matrix. However, its performance deteriorates when the connection matrix contains a large amount of noise. Inspired by the structure of the ideal clustering indication matrix, we propose regularizing the projection matrix to achieve more accurate clustering results. We investigate the effectiveness of bounded, positive, and sparse regularization terms.
There are several promising directions for future research in this field. Firstly, we can explore alternative approaches to enhance clustering by investigating different models, such as non-negative matrix factorization or semi-definite approximation techniques. Secondly, in addition to the regularization terms discussed before, other forms of regularization could also be explored to further improve performance. Given the complexity and challenges inherent in the community detection problem, numerous unexplored avenues remain, which we identify as potential areas for future research efforts. In future studies, we may explore nonlinear techniques, such as those proposed in [28], to further enhance clustering performance.